CARS SPECTRA OF COLD H_{2} CLUSTERS

\underline{KIRILL E. KUYANOV}, Department of Chemistry, University of Southern California, Los Angeles, CA 90089; MIKHAIL N. SLIPCHENKO, Department of Chemistry, Iowa State University, Ames, IA 50011; BOBBY KHALAJESTANI, RUSSELL SLITER and ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089.

Neat H_{2} clusters as well as H_{2} clusters embedded in helium droplets have been obtained in a pulsed cryogenic nozzle beam expansion and studied via Coherent anti-Stokes Raman Scattering (CARS) technique. The motivation for these experiments is to spectroscopically observe superfluidity in molecular hydrogen clusters. Results on large hydrogen clusters of about 10^{4} molecules presented. We have studied the composition and structure of the clusters obtained from the H_{2} gas with different concentrations of para-H_{2} and ortho-H_{2} molecules. CARS spectra indicate the enrichment in ortho-H_{2} in the clusters.